OpenStudy (anonymous):

Chester makes 1L of a 0.01M HCl solution and measure the pH. Burt then adds 5.80 g of sodium chloride to Chester’s solution and measures the pH. Calculate the difference between the two pH values obtained.

OpenStudy (photon336):

pH = -LOG[H+] This means that the pH is the -log of the [H+] which is the concentration of acid.

OpenStudy (photon336):

because HCl dissociates completely, we know that we'll have 0.01M of H+

OpenStudy (photon336):

@MASTERPARKOUR

OpenStudy (anonymous):

I understand that, but how do I calculate the change in pH after the addition of NaCl?

OpenStudy (photon336):

@aaronq @sweetburger

OpenStudy (aaronq):

There won't be any noticeable change in pH because the dissociation constant for HCl is really high, 1.3*10^6. Here the chloride ion would behave as a base, thus the increase in \([Cl^-]\) would cause the reformation of HCl, decreasing \([H^+]\) available, in the reaction: \(Cl^-+H^+\rightleftharpoons HCl\) The K for this reaction is the reciprocal of the dissociation constant for HCl. \(K=\dfrac{1}{1.3*10^6}=7.7*10^{-7}\) \(\sf K=\dfrac{[HCl]}{[H^+][Cl^-]}=\dfrac{[HCl]}{(0.01-[HCl])((0.01+0.1)-[HCl])}\) (Note: I skipped showing the I.C.E table for the result above) \(\sf [HCl]\approx8.5*10^{-10}M\) The amount of \([H^+]\) is 8 orders of magnitude lower, so it's very insignificant.

OpenStudy (aaronq):

The amount of \([H^+]\) taken up** i meant.